OpenStudy (anonymous):

A sequence of numbers is shown. 1, 3, 17, 115, ... Which describes how to find the next term? A. Add 14 to the previous term. B. Add 2 to the previous term. C. Multiply the previous term by 4, and then subtract 7 from the result. D. Multiply the previous term by 7, and then subtract 4 from the result.

OpenStudy (tkhunny):

ANY method one might select. There are INFINTELY many possiblities. The question should read, "This sequence was generated by one of the following methods. Select the one that works." A: 1+14 = 15 -- That one doesn't work. B: 1+2 = 3 -- Looking good. 3 + 2 = 5 -- Whoops! That's not 17. C: 4*1 - 7 = -3 -- That doesn';t work. Only one let! D: 7*1 - 4 = 3 -- Good. 7*3 - 4 = 17 -- Nice. 7*17 - 4 = 115 -- I think we're done.